# Fine-tuning Vision Language Models with Graph-based Knowledge for Explainable Medical Image Analysis

### v1

@@ Abstract
Accurate staging of Diabetic Retinopathy (DR) is essential for guiding timely interventions and preventing vision loss. However, current staging models are hardly interpretable, and most public datasets contain no clinical reasoning or interpretation beyond image-level labels. In this paper, we present a novel method that integrates graph representation learning with vision-language models (VLMs) to deliver explainable DR diagnosis. Our approach leverages optical coherence tomography angiography (OCTA) images by constructing biologically informed graphs that encode key retinal vascular features such as vessel morphology and spatial connectivity. A graph neural network (GNN) then performs DR staging while integrated gradients highlight critical nodes and edges and their individual features that drive the classification decisions. We collect this graph-based knowledge which attributes the model’s prediction to physiological structures and their characteristics. We then transform it into textual descriptions for VLMs. We perform instruction-tuning with these textual descriptions and the corresponding image to train a student VLM. This final agent can classify the disease and explain its decision in a human interpretable way solely based on a single image input. Experimental evaluations on both proprietary and public datasets demonstrate that our method not only improves classification accuracy but also offers more clinically interpretable results. An expert study further demonstrates that our method provides more accurate diagnostic explanations and paves the way for precise localization of pathologies in OCTA images.
Keywords:
Vision language models Graph learning DR OCTA.
@@ 1 Introduction
In this paper, we present a novel method for training a VLM agent that enhances the interpretability of DR staging by providing a textual interface which allows clinicians to directly interact with the DR staging model. We achieve this by first integrating a GNN to capture complex spatial relationships in OCTA images of the deep vascular complex (DVC) for DR staging, and then transforming this graph-based knowledge into structured, table-formatted texts for a teacher model to generate fine-tuning data. We subsequently fine-tune VLM models with direct image inputs. The results show enhanced performance in both classification and interpretation tasks, where only OCTA images and a few lines of background knowledge are provided as prompts, marking a promising step toward models that can both classify and explain their predictions in a clinically relevant manner.
@@ 2.3 Feature and Location Attribution
[…]
@@ 3 Experiments
We conduct comprehensive experiments with 4 state-of-the-art VLMs to evaluate our method’s quantitative and qualitative performance in DR staging and explanation. Code is available at https://anonymous.4open.science/r/GFT-DR/.
@@
Evaluation Metrics. For classification performance, we evaluate all models using balanced accuracy, precision, and recall rate. For explanation performance, we first follow previous works [11, 13] and use the ground-truth diagnosis and graph-based knowledge to prompt the teacher model to generate a set of responses as standards, and then ask it to compare with the candidate models’ responses and give scores (0-100) based on the quality of the explanations. Furthermore, we present 48 responses of each model to two ophthalmology experts. These 48 responses are generated by randomly shuffling the outputs from 6 models across 8 samples. The experts then rank and assign scores to the responses based on three criteria: overall accuracy, correct localizations, and helpfulness. The experts’ ratings for each model are then averaged to obtain the final scores.
@@ 4.1 Diabetic Retinopathy Staging Results
Table 1 presents the DR staging performance on our proprietary dataset. Our graph-knowledge-fine-tuned (GFT) models consistently outperform the baseline vision-language models (BS) and standard fine-tuning (FT) approaches. GFT-Llama 3.2 11b achieves a balanced accuracy of 0.678, only slightly lower than the specialized GNN’s performance, reported as 0.689 on the identical test set in [14]. Across architectures, GFT brings an average improvement of 14.8% in balanced accuracy to FT, demonstrating the effectiveness of integrating graph.
@@ 4.2 Quality of Explanation
Table 3 presents a quantitative evaluation of explanation quality across different model configurations. We assess explanation quality through both automated metrics (teacher model evaluation) and human expert review. The teacher model scores reflect alignment with ground truth explanations generated from the graph-based knowledge, while expert scores evaluate clinical relevance and accuracy. The Loc. metric quantifies a model’s ability to correctly identify specific retinal regions containing pathological features across all test cases.

### v2

@@ Abstract
Accurate staging of Diabetic Retinopathy (DR) is essential for guiding timely interventions and preventing vision loss. However, current staging models are hardly interpretable, and most public datasets contain no clinical reasoning or interpretation beyond image-level labels. In this paper, we present a novel method that integrates graph representation learning with vision-language models (VLMs) to deliver explainable DR diagnosis. Our approach leverages optical coherence tomography angiography (OCTA) images by constructing biologically informed graphs that encode key retinal vascular features such as vessel morphology and spatial connectivity. A graph neural network (GNN) then performs DR staging while integrated gradients highlight critical nodes and edges and their individual features that drive the classification decisions. We collect this graph-based knowledge which attributes the model’s prediction to physiological structures and their characteristics. We then transform this reasoning into textual descriptions for VLMs. We perform instruction-tuning with these textual descriptions and the corresponding image to train a student VLM. This final agent can classify the disease and explain its decision in a human interpretable way solely based on a single image input. Experimental evaluations on both proprietary and public datasets demonstrate that our method not only improves classification accuracy but also offers more clinically interpretable results. An expert study further demonstrates that our agent provides more accurate diagnostic explanations and enables precise localization of pathologies in OCTA images.
@@ 1 Introduction
In this paper, we present a novel method for training a VLM agent that enhances the interpretability of DR staging by providing a textual interface which allows clinicians to directly interact with the DR staging model. We achieve this by first integrating a GNN to capture complex spatial relationships in OCTA images of the deep vascular complex (DVC) for DR staging, and then transforming this graph-based knowledge into structured, table-formatted texts for a teacher model to generate fine-tuning data. We subsequently fine-tune VLM models with direct image inputs. The results show enhanced performance in both classification and interpretation tasks, where only OCTA images and a few lines of background knowledge are provided as prompts, marking a promising step toward models that can both classify and explain their predictions in a clinically relevant manner. In that, our method is related to current advances in reasoning language models.
@@ 2.3 Feature and Location Attribution
$\mathrm{IG}_{i}(\mathbf{x})=(x_{\text{input},i}-x_{\text{baseline},i})\times\int_{0}^{1}\frac{\partial f(\mathbf{x}_{\text{baseline}}+\alpha(\mathbf{x}_{\text{input}}-\mathbf{x}_{\text{baseline}}))}{\partial x_{i}}\,d\alpha,$ (1)
@@ 3 Experiments
We conduct comprehensive experiments with 4 state-of-the-art VLMs to evaluate our method’s quantitative and qualitative performance in DR staging and explanation. Code is available at https://github.com/chenjun-li/GFT.
@@
Evaluation Metrics. For classification performance, we evaluate all models using balanced accuracy, precision, and recall rate. For explanation performance, we first follow previous works [11, 13] and use the ground-truth diagnosis and graph-based knowledge to prompt the teacher model to generate a set of responses as standards, and then ask it to compare with the candidate models’ responses and give scores (0-100) based on the quality of the explanations. Furthermore, we present 48 responses of each model to two ophthalmology experts. These 48 responses are generated by randomly shuffling the outputs from 6 models across 8 samples. The experts then rank and assign scores to the responses based on three criteria: overall accuracy, correct localizations, and helpfulness. A quadrant-based system is used to verify localizations correspondence. A region is marked as correct only when both experts agree. The ratings for each model are then averaged to obtain the final scores.
@@ 4.1 Diabetic Retinopathy Staging Results
Table 1 presents the DR staging performance on our proprietary dataset. Our graph-knowledge-fine-tuned (GFT) models consistently outperform the baseline vision-language models (BS) and standard fine-tuning (FT) approaches. GFT-Llama 3.2 11b achieves a balanced accuracy of 0.678, only slightly lower than the specialized GNN’s performance, reported as 0.689 on the identical test set in [14]. Across architectures, GFT brings an average improvement of 14.8% in balanced accuracy to FT, demonstrating the effectiveness of integrating graph. The confidence intervals from the 5-fold cross-validation are: balanced accuracy $\pm 10.13\%$, precision $\pm 5.67\%$, and recall $\pm 4.47\%$, respectively.
@@ 4.2 Quality of Explanation
Table 3 presents a quantitative evaluation of explanation quality across different model configurations. We assess explanation quality through both automated metrics (teacher model evaluation) and human expert review. The teacher model scores reflect alignment with ground truth explanations generated from the graph-based knowledge, while expert scores evaluate clinical relevance and accuracy. The Loc. metric quantifies a model’s ability to correctly identify specific retinal regions containing pathological features across all test cases. The inter-rater weighted agreement is $\kappa=0.83$, indicating almost perfect agreement.
@@ 5 Conclusion
{credits}
[…]
The authors have no competing interests to declare that are relevant to the content of this article.